OpenStudy (anonymous):

how do you simplify -3radical 9

OpenStudy (anonymous):

\(\sqrt{9}=3\)

OpenStudy (anonymous):

so \(-3\sqrt9=-3\times 3\)

OpenStudy (jdoe0001):

\(\bf -3\sqrt{9}\quad ?\)